Minutes of the code management group meeting 2016-09-23

From UKCA
Jump to: navigation, search

Code management group minutes


Present

Met Office: John Hemmings, Fiona O'Connor, Nick Savage, Mohit Dalvi, Jane Mulcahy, Colin Johnson, Alan Hewitt

Cambridge: Luke Abraham, Alex Archibald, Paul Griffiths

Leeds: Mark Richardson

Reading: Nicolas Belloiun

Apologies:

Actions from this meeting

One off actions

  1. (ongoing) GM to check 8.4 versus 7.3 (and possibly vn8.4 vs 10.x) GLOMAP-mode code and report any missing bug fixes on the bugfix page.
  2. (ongoing) FOC to incorporate Bug #35 and others into dry deposition work
  3. (ongoing) LA to migrate PUMA UM tickets to MOSRS
  4. (ongoing) UM #368 (RAFEPS value). Error not encountered at UM10.x but LA to consult Oliver Wild regarding background and implications.
  5. (ongoing) LA to check doc for NVOC factor and open ticket to remove it from code
  6. (ongoing) CJ to lodge STASHmaster changes to diagnostics that should make use of UKCA time availability codes
  7. (ongoing) CJ,FOC,AH,AS,LA and John H to decide best way forward for implementing GLOMAP-climatology code and documentation.
  8. (ongoing) JM,BJ,NB to discuss incorporation of Nicolas' RADAER description into UKCA (or separate UMDP). JM to take lead.
  9. (ongoing) FOC to open ticket + Bugfixes Table entry related to 'Unable to calculate tropopause' warnings
  10. (updated) MD to setup Strat+Trop suite as soon as GA7.1 is frozen and also add to Rose-stem. MD set-up/replace current Rose stem job with GA7.1 jobs.
  11. (ongoing) CJ to involve NS in discussions with Dynamics and UM-sys teams regarding movement of UKCA call, for effective implementation of water vapour feedback.
  12. (ongoing) FOC to recommend at next Exec meeting a policy of allowing only CMG to (certify and) distribute any UKCA configurations for science purpose.
  13. (new) MD to chase CMS about MASS archiving
  14. (new) NS to collate all code that could be removed and put on Wiki
  15. (new) LA to do doodle poll for end of Nov/ start of Dec

Standing actions

  1. (ongoing - standing item) CMG to monitor list of bugs and circulate between group for owners.
  2. (ongoing - standing item) At CMG future meetings there should be included an appendix listing the open bugs. If a ticket is included it is not an open bug (as the owner of ticket is person doing lodging).
  3. (ongoing - standing item) JM to continue to link/send powerpoint results (re: GA7.0/GA7.1), as new configurations released and link/put files on UKCA wiki.
  4. (ongoing - standing item) ALL with access to keep monitoring ukca_bugs email list.
  5. (ongoing - standing item) ALL to think about if code changes will change KGO and let systems team know in good time. This should be a standard part of our working practice.
  6. (ongoing - standing item) ALL to review and remove obsolete code

Teleconference Numbers

  • UK Freefone:0800 9171956
  • Participant passcode: 89264837 then #

Agenda

  1. Agree agenda
  2. Actions from the last meeting (2016-07-22) including standing items
  3. Targets for UMvn10.6 code freeze
  4. Status of GA7.1
  5. Rose suites at UMvn10.4/10.5 & Progress towards GA7.1 & StratTrop configuration
  6. Bugs/bugfixes
  7. Problems with N-R solver
  8. Update on GungoHo!/LFRic
  9. AOB/DNM
  10. Date of next meeting


Minutes

Actions from 2016-07-22

  1. (ongoing) GM to check 8.4 versus 7.3 (and possibly vn8.4 vs 10.x) GLOMAP-mode code and report any missing bug fixes on the bugfix page.
    • MR to remind GM to check against 10.x
  2. (ongoing) GLOMAP-mode coupling to FJX. AA to see if Paul Griffiths can be asked to do this.
    • done. This is ongoing work. PG standalone FJX working but no progress on coupling.PG to open/check for ticket to do work.
  3. (ongoing) FOC to incorporate Bug #35 and others into dry deposition work
    • To be done on work funded through LSTM project, which will add flexibility. Bug will remain open currently.
  4. (ongoing) FOC/LA to find someone to co-ordinate migration of PUMA UM tickets to MOSRS
    • (ongoing) LA to do
  5. UM #368 (RAFEPS value). Error not encountered at UM10.x but LA to consult Oliver Wild regarding background and implications.
    • not done. CJ to add to this. See later
  6. (ongoing) Bug #38: MD to lodge the switch for controlling addition of SO2 from explosive volcanoes.
    • done. In code review
  7. (ongoing) LA to check doc for NVOC factor and open ticket to remove it from code
    • ongoing
  8. CJ to lodge STASHmaster changes to diagnostics that should make use of UKCA time availability codes
    • ongoing. working on time availability for UKCA diagnostics on UKCA timestep. Update to STASHmaster for next code freeze
  9. FOC to supervise work by Joao for UM ticket #987 (Checks in UKCA iBVOC emissions to prevent duplication of online and offline emissions)
    • done. Gerd has taken over ticket as a fix had to go in. Currently in SciTech review
  10. (Ongoing)NS to keep Bug #58 (Blocking Mode 3 in FastJX) ticket open but low priority.
    • will remain on bugs table. Remove as action
  11. NS to update MR on problems encountered due to module-defined variables and OpeNMP in Convection, with possible implications for GLOMAP OpenMP work
    • done. MR starting work
  12. CJ,FOC,AH,AS,LA and John H to decide best way forward for implementing GLOMAP-climatology code and documentation.
    • Leave as ongoing. Would like to avoid duplication for bugfix purposes. Discuss at a later meeting
  13. JM,BJ,NB to discuss incorporation of Nicolas' RADAER description into UKCA (or separate UMDP).
    • The same might apply for ACTIVATE & GLOMAP climatology. It might be best to incorporate into one for GLOMAP-mode Aerosol cloud- & aerosol-radiation interactions. There may also be links with CASSIM scheme? Currently the aerosol climatology (s54) is documented in the Reconfiguration documenation
    • ongoing. JM to take lead
  14. FOC to talk to Paul S. regarding possibility of lodging Mark's basic (hardwired) column-based/ OpenMP approach at UM10.5 and more flexibility in follow-on ticket. NS/Alistair can follow-up.
    • done
  15. Bug #72: NS to take a look at Dry deposition/ surface resistance warnings and explore possibility of a quick work-around.
    • done, ongoing
    • The problem with this scheme as it is implemented is that if something is missing, the model carries on. Some species that have 2D (old) dry-deposition values are missing one or other surface resistance or diffusion values needed by the interactive scheme.
    • NS cannot do any more work. One possible solution would be to include this in the current dry deposition work. We would need to try setting them to be not depositied, and see impacts on kgo. Warnings should be turned to an error at the same time.
  16. FOC to open ticket + Bugfixes Table entry related to 'Unable to calculate tropopause' warnings
    • ongoing. separate to #74, but will incorporate work incorporate to this ticket
  17. MD to setup Strat+Trop suite as soon as GA7.1 is frozen and also add to Rose-stem.
    • ongoing
    • GA7.1 is available, although some things still need to be done. Release job at 10.6 GA7.1
  18. FOC,JM, others to evaluate above job and ALL to decide about recommending this as the next UKCA Release job.
    • done - see above
  19. JM to enquire with David W. regarding addition of GA7.1 to Rose-stem.
    • There is a ga7_dev job already, which will contain most GA7.1 changes. There will also be a separate UKESM rose stem job, but doesn't include GA7.1 changes
    • It was decided to replace current eg_strattrop with ga7.1_strattrop
    • ACTION: MD to set-up/replace current job
  20. CJ to involve NS in discussions with Dynamics and UM-sys teams regarding movement of UKCA call, for effective implementation of water vapour feedback.
    • ongoing
  21. FOC to recommend at next Exec meeting a policy of allowing only CMG to (certify and) distribute any UKCA configurations for science purpose.
    • ongoing
  22. FOC to set-up Doodle poll for next CMG (possibly for week of 12th September).
    • done

Standing items:

  1. (ongoing - standing item) CMG to monitor list of bugs and circulate between group for owners.
  2. (ongoing - standing item) At CMG future meetings there should be included an appendix listing the open bugs. If a ticket is included it is not an open bug (as the owner of ticket is person doing lodging).
  3. (ongoing - standing item) JM to continue to link/send powerpoint results (re: GA7.0), as new configurations released and link/put files on UKCA wiki.
  4. (ongoing - standing item) ALL with access to keep monitoring ukca_bugs email list.
  5. (ongoing - standing item) ALL to think about if code changes will change KGO and let systems team know in good time. This should be a standard part of our working practice.

Targets for UMvn10.6 code freeze

UMvn10.6 tickets

There will possibly be an intermediate milestone between 10.6 & 10.7: 10.6.1

  • #2036 - CJ (2yr run) to run eval suite on his run and run from Mohit (20yrs). 2nd year of runs. JM to run maverick as well. CJ to run on for longer to run tests next week.
  • LA to check MD branch, possibly SciTech review.
  • CJ might not get time profile stuff ready in time. NS to SciTech it.

Status of GA7.1

See some updates above. GA7.1 has been frozen (mid August). DMS emissions have been scaled, which brings ERF back in line. However, there is a missing a source of marine aerosol, and more marine organic aerosol has been added as a temporary fix. The DMS scaling is not for StratTrop. This is documented under a GMED ticket (#275).

Rose suites at UMvn10.4/10.5 & Progress towards GA7.1 & StratTrop configuration

MD porting one over to MONSooN.

  • ACTION: MD to chase CMS about MASS archiving

Would need to email CMS for ARCHER port once its on MONSooN.

  • FOC to follow-up with AS

Data available on JASMIN

Bugs/Bugfixes

Open Bugs

# Description Ticket Owner Status/ Updates Comments
31 FP exceptions in asad_sparse_vars
or asad_spimjmp
#368 (MOSRS) Nick Savage leave open. vn10.x seems to be better
35 Error in labelling: Surface resistance <--> stomatal resistance ? MOSRS #1409 Fiona O'Connor open - folded into DD. FOC to update
38 Missing logical for volcanic so2 emissions. MOSRS #1350 Mohit Dalvi Code Review soon to be closed
50 ukca_solve_coagnucl failures #319 (MOSRS) Mohit Dalvi/Graham Mann? Not encountered any more ? dissapeared at vn10.x
57 Incorrect logic for DMS anciliaries required #756 (MOSRS) Nick Savage trivial. Might close as wontfix, but might want to do in RAQ
58 Fastj-X runs on patches all in dark #1360 (MOSRS) Nick Savage don't think as important, not a bug. Remove from list, but keep ticket. NS to do
68 HO2 heterogeneous loss rate 1st order(?) No ticket Nick Savage NS to look at that. Leave open to review
72 UKCA Dry Deposition: Surface resistance values
not set for some species
#1250 Alan Hewitt? discussed above, have way forward
74 Bug in ukca_tropopause with definition of PV tropopause #1283 Steven Hardiman Change ownership? discussed above, wrap together issues. Fix is separate from #75
75 Bug in STE fluxes #1367 Fiona O'Connor could be endgame. Needs further work
80 H2O feedback broken under ENDGAME UM#1646 Colin Johnson discussed above, keep open
81 Runs with segment/ restart interval < 5 days
fail due to NetCDF emiss update bug
UM#948
being lodged under UM#970
Alistair Sellar AS lodging this, ancil ref time. Will be lodged
83 error in the indexing of the tot_tracer array in the
convective plume scavenging
UM#2036 Nick Savage discussed above
85 Discontinuity in photolysis rate diagnostics UM#2286 Fiona O'Connor mentioned, FOC to investigate. Diagnostics to still go in

Problems with N-R solver

CJ is not sure where to start, and is looking at MR's problems. Getting lots of

convergence not achieved in halving step

messages. The issues are much worse with a coupled ocean, and problems are seen near the South Pole. The 2nd month performance is worse. When debugging, production terms are seen for species which shouldn't have production terms. DMS, C5H8 etc.

FOC suggests to look at before & after issues, and also suggests having a special meeting to discuss the solver.

  • CJ to work on this. Leave open. Will send some more stuff to LA/round

Update on GungHo!/LFRic

Not much progress. Have a secondment, but is running training currently. Should get time in October. Currently no transport at all. The first tracer tests were to be ENDGame anyway

AOB

Code

New standing item:

  • ACTION (standing item): ALL to review and remove obsolete code
  • ACTION: NS to collate all code that could be removed and put on Wiki

Date of next meeting

  • ACTION: LA to do doodle poll for end of Nov/ start of Dec.